The following is the information from the national income accounts for a hypothetical country: GDP Rs. 6000.00 Gross Investment 800.00 Net Investment 200.00 Consumption 4000.00 Govt. purchases of goods & services 1100.00 Govt. Budget Surplus 30.00.

What is aggregate Saving ??


1
Expert's answer
2021-04-12T08:48:07-0400

GDP: Y = Rs. 6000

Gross Investment I = 800

Net Investment NI = 200

Consumption C = 4000

Govt. purchases of goods & services G = 1100

Govt. Budget Surplus BS = T - G = 30


a) Depreciation = I - NI = 800 - 200 = 600.

NDP = Y - depreciation = 6000 - 600 = 5400.


b) Net exports NX = Y - C - I - G = 6000 - 4000 - 800 - 1100 = 100.


c) Govt. taxes minus transfers T = G + BS = 1100 + 30 = 1130.


c) Disposable personal income Yd = Y - T = 6000 - 1130 = 4870.


e) Saving S = Yd - C = 4870 - 4000 = 870.